# Estimating Model Performance Under Covariate Shift Without Labels

### v2

@@
Estimating Model Performance Under Covariate Shift Without Labels
Jakub Białek11footnotemark: 122footnotemark: 2
NannyML NV
Interleuvenlaan 62, Belgium
3001 Heverlee
jakub@nannyml.com
&Wojtek Kuberski 11footnotemark: 1
NannyML NV
Interleuvenlaan 62, Belgium
3001 Heverlee
wojtek@nannyml.com
&Nikolaos Perrakis 11footnotemark: 1
NannyML NV
Interleuvenlaan 62, Belgium
3001 Heverlee
nikos@nannyml.com &Albert Bifet
AI Institute, University of Waikato
Hamilton, New Zealand
LTCI, Telecom Paris,
IP Paris, France
abifet@waikato.ac.nz
Machine learning models often experience performance degradation post-deployment due to shifts in data distribution. It is challenging to assess post-deployment performance accurately when labels are missing or delayed. Existing proxy methods, such as drift detection, fail to measure the effects of these shifts adequately. To address this, we introduce a new method for evaluating classification models on unlabeled data that accurately quantifies the impact of covariate shift on model performance and call it Probabilistic Adaptive Performance Estimation ($\mathrm{PAPE}$). It is model and data-type agnostic and works for any performance metric. Crucially, $\mathrm{PAPE}$ operates independently of the original model, relying only on its predictions and probability estimates, and does not need any assumptions about the nature of the shift, learning directly from data instead. We tested $\mathrm{PAPE}$ using over 900 dataset-model combinations from US census data, assessing its performance against several benchmarks through various metrics. Our findings show that $\mathrm{PAPE}$ outperforms other methodologies, making it a superior choice for estimating the performance of classification models.
@@ 1 Introduction
The final step in the machine learning (ML) model development process is to evaluate the model on an unseen dataset known as test set. Successful performance on this dataset, indicating potential business value, typically leads to the model’s deployment to production. The expectation is that the model’s performance on the test set will be replicated with production data [13]. Yet, this assumption often fails due to data distribution shifts that lead to model performance degradation [19]. Therefore, ongoing monitoring of the model’s performance is essential to ensure it continues to meet expected outcomes. However, in many scenarios, real-time performance assessment is challenging because true labels are either unavailable or significantly delayed.
@@
Accurate performance estimates without labels can mitigate potential losses. With these estimates, the impact of changes in model performance on business operations can be quantified, allowing for informed decisions. These may include altering how model predictions are utilized in downstream processes, discontinuing their use altogether, or determining whether model retraining is required.
@@
  • We introduce a novel algorithm called $\mathrm{PAPE}$ (Probabilistic Adaptive Performance Estimation) for assessing model performance under covariate shift. $\mathrm{PAPE}$ enables estimation of various classification performance metrics and is designed to be both model and data-type agnostic. Importantly, it autonomously learns from data without requiring user input regarding the nature of the covariate shift.
  • We present both theoretical and empirical analyses demonstrating how $\mathrm{PAPE}$ can effectively yield accurate performance estimations.
  • We develop a comprehensive evaluation framework specifically tailored for assessing performance estimation methods. This framework includes datasets, their preprocessing steps, implementation of benchmarks suitable for tabular data, performance metrics of interest, and innovative evaluation metrics relevant for model monitoring context.
  • Our extensive evaluations reveal that $\mathrm{PAPE}$ significantly surpasses all existing benchmarks across all metrics evaluated, showing its superior performance estimation capabilities.
The paper is organized as follows. In Section 2, we present the related work. In Section 3, we introduce the $\mathrm{PAPE}$ algorithm, presenting its practical implementation, theoretical foundations, guarantees, and limitations. Section 4 describes the experimental methodology we used to assess performance estimation methods, and presents the results. Section 5 summarizes the results.
@@ 2 Related work
Importance Weighting (IW) [9] is a simple but powerful method of estimating model performance. IW weighs the labeled test data with a likelihood ratio of observing such input production data distribution and calculates a weighted performance metric. Extended version of IW [1] takes into account the nature of covariate shift and its potential impact on performance, as defined by a user. Confidence-based methods form another family of approaches. Average threshold confidence (ATC) [6] learns a threshold on probability estimates and predicts accuracy using the fraction of predictions that exceed the threshold. Difference of confidence (DoC) [8] uses the difference between probability estimates from test to production data and fits a regression model that maps this difference into accuracy. DoC is an example of a distance-based approach that measures the distance between test and production data and identifies the relationship between the distance and the expected performance. Another example is using Fréchet distance and linear regression [3]. Other approaches, such as Contrastive Automated Model Evaluation (CAME) [16], use contrastive learning. Finally, reverse testing (RT) [5] methods use the monitored model predictions as labels to train a reverse model on production data. Then, they assume that the reverse model’s performance on the test dataset indicates the monitored model’s performance on the production data.
3 Probabilistic Adaptive Performance Estimation ($\mathrm{PAPE}$)
In this section, we present $\mathrm{PAPE}$ (Probabilistic Adaptive Performance Estimation), a new algorithm for assessing model performance under covariate shift. In Section 3.1, we provide a raw version of the $\mathrm{PAPE}$ algorithm. Section 3.2 gives a verbose description of the notation and provides a proof for using the main formula in Algorithm 1. In Section 3.3, we describe the practical limitations of the new algorithm.
@@ 3.1 The algorithm
The steps, required inputs, and outputs of the $\mathrm{PAPE}$ algorithm are detailed in Algorithm 1.
Algorithm 1 Probabilistic Adaptive Performance Estimation ($\mathrm{PAPE}$)
@@
  • Labelled reference data: inputs $x_{ref}$, labels $y_{ref}$ and number of observations $n_{ref}$
  • Unlabelled production data: inputs $x_{prod}$ ; $n_{prod}$ observations
@@
  • Actionable predictions of $f$ on production data: $g(f(x_{prod}))$
  • Function $m$ that computes the performance metric: $m\left(g(f(x)),y\right)$
@@
  • Estimation of performance metric $m$ of classifier $f$ on unlabelled production data, that is $\widehat{m}_{prod}$
@@
    Use $c$ to get calibrated probabilities on production data: $c\left(f(x_{prod})\right)$
@@
    $\widehat{m}_{prod}=\mathrm{mean}_{x\sim prod}\left[c\left(f(x)\right)m\left(g(f(x)),1\right)+\left(1-c\left(f(x)\right)\right)m\left(g(f(x)),0\right)\right]$ (1)
@@ 3.2 Theoretical guarantees
Fix a distribution $D$. Consider an uncalibrated classifier $f$ that for each observation $x\in X$ calculates probability estimates (referred to as scores, $s$) $f:X\rightarrow S$, where $S=[0,1]$. Assume that there is a function $c$ that calibrates scores $s$ into probabilities, that is $P[y=1|f(x)=s]=c(s)$ for all $s\in S$. Suppose there is a function $g$ that maps scores returned by the classifier into actionable predictions. Consider a performance metric of interest $m$ that is calculated based on predictions of the classifier $g(f(x))$ and corresponding true labels $y$ that is $m\left(g(f(x)),y\right)$.
Lemma.
@@
$\widehat{m}_{(f,g,D)}=\mathrm{E}_{D}\left[c(f(x))m\left(g(f(x)),1\right)+\left(1-c(f(x))\right)m\left(g(f(x)),0\right)\right]$ (2)
@@
Since all the expected values and probabilities are calculated over distribution $D$ we will omit the subscript $D$ for better readability. We will additionally consider that $c$ calibrates scores $s$ only approximately, with expected absolute calibration error $\varepsilon$, that is:
$E[y|f(x)=s]=P\left[y=1|f(x)=s\right]=c\left(s\right)+\epsilon_{s}$ (3)
$\sum_{s\in S}P[f(x)=s]E[|y-c(s)|f(x)=s|]=\sum_{s\in S}P[f(x)=s]\left|\epsilon_{s}\right|=\varepsilon$ (4)
@@
$m_{\left(f,g\right)}$ $=\mathrm{E}\left[m\left(g(f(x)),y\right)\right]$
$=\sum_{s\in S}P[f(x)=s]\mathrm{E}[m(g(f(x)),y)|f(x)=s]$
$=\sum_{s\in S}P[f(x)=s]\mathrm{E}[m(g(s),1)P[y=1|f(x)=s]+m(g(s),0)P[y=0|f(x)=s]]$
$=\sum_{s\in S}P[f(x)=s]\mathrm{E}[m(g(s),1)(c(s)+\epsilon_{s})+m(g(s),0)(1-(c(s)+\epsilon_{s}))]$
$=\sum_{s\in S}P[f(x)=s]\mathrm{E}[m(g(s),1)c(s)+m(g(s),0)(1-c(s))]$
$+\sum_{s\in S}P[f(x)=s]\mathrm{E}[\epsilon_{s}(m(g(s),1)-m(g(s),0))]$
$=\widehat{m}_{(f,g)}+\sum_{s\in S}P[f(x)=s]\mathrm{E}[\epsilon_{s}(m(g(s),1)-m(g(s),0))]$
If calibration is perfect, that is $\epsilon_{s}=0,\forall s\in S$, then the estimated metric equals the true metric. In other cases, performance estimation is as accurate as the calibration process. For example, for performance metrics for which $\max[m(g(s),1)-m(g(s),0)]=1$ (like accuracy, confusion matrix elements etc.) the absolute error of the performance metric estimation is limited by the expected absolute calibration error:
@@
To use equation (2) for estimating the performance of $f$ on unlabelled data from production distribution $D_{prod}$ we need $c$ a regressor that calibrates the scores on $D_{prod}$. We only have labels for the reference distribution to train the calibrator $c$ However, when probabilities are calibrated under reference distribution $D_{ref}$ and with respect to density rates $w(x)_{ref\rightarrow prod}=\frac{p(x)_{prod}}{p(x)_{ref}}$ they become calibrated under production distribution [11]. That is true under the assumption that the true conditional probabilities of labels given the inputs remain the same for reference and production distribution, that is - under covariate shift: $p(y|X)_{ref}=p(y|X)_{prod}$. The input distribution may still change so that $p(x)_{ref}\neq p(x)_{prod}$. The probability density rates $w(x)_{ref\rightarrow prod}$ describe the input distribution shift from reference to production. Calibrator $c$ provides calibrated probabilities under $D_{prod}$ when:
$\mathrm{E}_{D_{ref}}\left[w(x)_{ref\rightarrow prod}\left(y-c(f(x))\right)|f(x)=s\right]=0,\forall s\in S$ (6)
In steps 2 and 3 of Algorithm 1, we approximate weighted calibration under reference distribution and weights $w$ by fitting $w$-weighted regressor on reference data. The weights $w$ are estimated in steps 1a-1e with DRE as described in [14, chapter 2.7.5].
@@ 3.3 Limitations
Our method will not work if there is concept shift, that is - the conditional distribution of label given inputs changes between reference and production data $p(y|X)_{ref}\neq p(y|X)_{prod}$. However, even if that is not the case, the data may shift to a region that is not part of the reference distribution. That is, there is some region $X_{r}$ in the input space $X$ for which $p(X_{r})_{prod}>0$ but $p(X_{r})_{ref}=0$. In that case weighted calibration will not be possible, as all the weights $w$ will be equal to zero (that is there are no observations in reference data similar to the ones from production). Other concerns are the same as for every other method that learns from data and are mainly related with finite sample size (of either reference or production data). $\mathrm{PAPE}$ uses equation 1 that assumes that the probabilities are perfectly calibrated, which is rarely the case when working on finite-size data (due to both - difficulties in learning correct $w$ weights and the calibration process itself). Finally, even with perfectly calibrated probabilities, the realized sample performance may end up different from the estimated one due to uncertainty related to probabilities realizing into binary targets. Addressing these issues is a subject of our current research.
4 Experimental evaluation
This section describes our experimental setting that evaluates and compares the proposed method with the existing benchmarks. We describe the datasets we used in 4.1, how we created the ML models whose performance was estimated, and the experimental setup used for performance estimation evaluation in 4.2. Additionally, we provide practical implementation details of the evaluated benchmarks in 4.3.
@@ 4.1 Datasets
The datasets we use to evaluate the method come from Folktables [4]. Folktables uses US census data and preprocesses them to create a set of binary classification problems. We use the following tasks: ACSIncome, ACSPublicCoverage, ACSMobility, ACSEmployment, ACSTravelTime. For each of the five prediction tasks listed above, we fetch Folktables data for all 50 US states spanning four consecutive years (2015-2018). This gives us 250 datasets.

### v3

@@
Abstract
Machine learning models often experience performance degradation post-deployment due to shifts in data distribution. It is challenging to assess model’s performance accurately when labels are missing or delayed. Existing proxy methods, such as drift detection, fail to measure the effects of these shifts adequately. To address this, we introduce a new method, Probabilistic Adaptive Performance Estimation (PAPE), for evaluating classification models on unlabeled data that accurately quantifies the impact of covariate shift on model performance. It is model and data-type agnostic and works for various performance metrics. Crucially, PAPE operates independently of the original model, relying only on its predictions and probability estimates, and does not need any assumptions about the nature of the covariate shift, learning directly from data instead. We tested PAPE on tabular data using over 900 dataset-model combinations created from US census data, assessing its performance against multiple benchmarks. Overall, PAPE provided more accurate performance estimates than other evaluated methodologies.
@@ 1 Introduction
The final step in the machine learning (ML) model development process is to evaluate the model on an unseen dataset known as a test set. Successful performance on this dataset, indicating potential business value, typically leads to the model’s deployment to production. The expectation is that the model’s performance on the test set will be replicated with production data [1]. Yet, this assumption often fails due to data distribution shifts that lead to model performance degradation [2]. Therefore, ongoing monitoring of the model’s performance is essential to ensure it continues to meet expected outcomes. However, real-time performance assessment is challenging in many scenarios because true labels are either unavailable or significantly delayed.
@@
Accurate performance estimates without labels can address this risk. With these estimates, the impact of changes in model performance on business operations can be quantified, allowing for informed decisions. These may include altering how model predictions are utilized in downstream processes, discontinuing their use altogether, or determining whether model retraining is required.
@@
  • We introduce a novel algorithm called PAPE (Probabilistic Adaptive Performance Estimation) for assessing model performance without labels under covariate shift. PAPE enables estimation of various classification performance metrics and is designed to be both model and data-type agnostic. Importantly, it autonomously learns from data without requiring user input regarding the nature of the covariate shift.
  • We present theoretical and empirical analyses demonstrating how PAPE can effectively yield accurate performance estimations.
  • We develop an evaluation methodology specifically tailored to assess performance estimation methods. This includes a selection of relevant datasets, their preprocessing steps, an implementation of benchmarks suitable for tabular data and performance metrics of interest, and new evaluation metrics relevant to the model monitoring context.
  • Our extensive evaluations show that PAPE significantly outperforms existing benchmarks across datasets and metrics used.
In Section 2, we present the related work. In Section 3, we introduce the PAPE algorithm, presenting its implementation, theoretical foundations, guarantees, and limitations. Section 4 describes the evaluation methodology we used to assess performance estimation methods and presents the results. Section 5 summarizes the results and outlines future research directions.
@@ 2 Related work
Importance Weighting (IW) [6] is a simple but powerful method of estimating model performance. IW calculates a weighted performance metric by weighing the labeled reference data based on the likelihood ratio of observing them in the production distribution compared to the reference distribution. An extended version of IW [7] takes into account the nature of covariate shift and its potential impact on performance, as defined by a user. Confidence-based methods form another family of approaches. Confidence-based performance estimation (CBPE) [8] calibrates probability estimates using labeled reference data and calculates expected confusion matrix elements on production data to turn them into the desired performance metric. Average threshold confidence (ATC) [9] learns a threshold on confidence estimates and predicts accuracy using the fraction of predictions that exceed the threshold. Difference of confidence (DoC) [10] uses the difference between confidence estimates from test to production data and fits a regression model that maps this difference into the performance metric. DoC is an example of a distance-based approach that measures the distance between test and production data and identifies the relationship between the distance and the expected performance. Another example of this approach is using Fréchet distance and linear regression [11]. Other approaches, such as Contrastive Automated Model Evaluation (CAME) [12], use contrastive learning. Finally, reverse testing (RT) [13] methods use the monitored model predictions as labels to train a reverse model on production data. Then, they assume that the reverse model’s performance on the labeled reference dataset corresponds to the monitored model’s performance on the production data.
3 Probabilistic Adaptive Performance Estimation (PAPE)
In this section, we present PAPE algorithm. In  3.1, we describe the algorithm’s inputs, outputs, and steps. Section 3.2 gives a verbose description of the notation and provides proof for using the main formula in Algorithm 1.
@@ 3.1 The algorithm
The steps, required inputs, and outputs of the PAPE algorithm are detailed in Algorithm 1.
Algorithm 1 Probabilistic Adaptive Performance Estimation (PAPE)
@@
  • Labeled reference data: inputs $x_{ref}$, labels $y_{ref}$ and number of observations $n_{ref}$
  • Unlabeled production data: inputs $x_{prod}$ ; $n_{prod}$ observations
@@
  • Actionable predictions of $f$ on production data: $g\bigl{(}f(x_{prod})\bigr{)}$
  • Function $m$ that computes the performance metric: $m\bigl{(}g(f(x)),y\bigr{)}$
@@
  • Estimation of performance metric $m$ of classifier $f$ on unlabeled production data - $\widehat{m}_{prod}$
@@
    Use $c$ to get calibrated probabilities on production data: $c\bigl{(}f(x_{prod})\bigr{)}$
@@
    $\widehat{m}_{prod}=\mathrm{mean}_{x\sim prod}\left[c\bigl{(}f(x)\bigr{)}m\bigl{(}g(f(x)),1\bigr{)}+\bigl{(}1-c(f(x))\bigr{)}m\bigl{(}g(f(x)),0\bigr{)}\right]$ (1)
@@ 3.2 Theoretical guarantees
Consider a binary classification task with inputs $x\in X$ and targets $y\in Y$, where $Y=\{0,1\}$. Fix a distribution $D$ and assume that $(x,y)\sim D$. Assume an uncalibrated classifier $f$ that for each observation $x\in X$ calculates probability estimates $f(x)=\hat{P}[y=1|x]$, referred to as scores $s$, so that $f:X\rightarrow S$, where $S=[0,1]$. Suppose that there is a function $c$ that calibrates scores $s$ into probabilities, that is $c(s)=P[y=1|f(x)=s]$ for all $s\in S$. Assume a function $g$ that maps scores returned by the classifier into actionable predictions. Consider a performance metric of interest $m$ that is calculated based on predictions of the classifier $g(f(x))$ and corresponding true labels $y$ that is $m\bigl{(}g(f(x)),y\bigr{)}$.
Lemma 1.
@@
$\widehat{m}_{(f,g,D)}=\mathrm{E}_{D}\left[c\bigl{(}f(x)\bigr{)}m\bigl{(}g(f(x)),1\bigr{)}+\bigl{(}1-c(f(x))\bigr{)}m\bigl{(}g(f(x)),0\bigr{)}\right]$ (2)
@@
Since all the expected values and probabilities are calculated over distribution $D$ we will omit the subscript $D$ for readability. We will additionally consider that $c$ calibrates scores $s$ only approximately, with expected absolute calibration error $\varepsilon$, that is:
$\mathrm{E}\bigl{[}y|f(x)=s\bigr{]}=P\bigl{[}y=1|f(x)=s\bigr{]}=c(s)+\epsilon_{s}$ (3)
$\sum_{s\in S}P\bigl{[}f(x)=s\bigr{]}E\Bigl{[}\bigl{|}y-c(s)|f(x)=s\bigr{|}\Bigr{]}=\sum_{s\in S}P\bigl{[}f(x)=s\bigr{]}\left|\epsilon_{s}\right|=\varepsilon$ (4)
@@
$m_{\left(f,g\right)}$ $=\mathrm{E}\left[m\bigl{(}g(f(x)),y\bigr{)}\right]$
$=\sum_{s\in S}P\bigl{[}f(x)=s\bigr{]}\mathrm{E}\Bigl{[}m\bigl{(}g(f(x)),y\bigr{)}|f(x)=s\Bigr{]}$
$=\sum_{s\in S}P\bigl{[}f(x)=s\bigr{]}\mathrm{E}\Bigl{[}m\bigl{(}g(s),1\bigr{)}P\bigl{[}y=1|f(x)=s\bigr{]}+m\bigl{(}g(s),0\bigr{)}P\bigl{[}y=0|f(x)=s\bigl{]}\Bigr{]}$
$=\sum_{s\in S}P\bigl{[}f(x)=s\bigr{]}\mathrm{E}\Bigl{[}m\bigl{(}g(s),1\bigr{)}\bigl{(}c(s)+\epsilon_{s}\bigr{)}+m\bigl{(}g(s),0\bigr{)}\Bigl{(}1-\bigl{(}c(s)+\epsilon_{s}\bigr{)}\Bigr{)}\Bigr{]}$
$=\sum_{s\in S}P\bigl{[}f(x)=s\bigr{]}\mathrm{E}\Bigl{[}m\bigl{(}g(s),1\bigr{)}c(s)+m\bigl{(}g(s),0\bigr{)}\bigl{(}1-c(s)\bigr{)}\Bigr{]}$
$+\sum_{s\in S}P\bigl{[}f(x)=s\bigr{]}\mathrm{E}\Bigl{[}\epsilon_{s}\Bigl{(}m\bigl{(}g(s),1\bigr{)}-m\bigl{(}g(s),0\bigr{)}\Bigr{)}\Bigr{]}$
$=\widehat{m}_{(f,g)}+\sum_{s\in S}P\bigl{[}f(x)=s\bigr{]}\mathrm{E}\Bigl{[}\epsilon_{s}\Bigl{(}m\bigl{(}g(s),1\bigr{)}-m\bigl{(}g(s),0\bigr{)}\Bigr{)}\Bigr{]}$
If calibration is perfect, that is $\epsilon_{s}=0,\forall s\in S$, then the estimated metric equals the true metric. ∎
In other cases, performance estimation is as accurate as the calibration process. For example, for performance metrics for which $\max[m(g(s),1)-m(g(s),0)]=1$ the absolute error of the performance metric estimation is limited by the expected absolute calibration error:
@@
To use equation (2) for estimating the performance of $f$ on unlabeled data from production distribution $D_{prod}$ we need $c$ a regressor that calibrates the scores on $D_{prod}$. We only have labels for the reference distribution to train the calibrator $c$ However, when probabilities are calibrated under reference distribution $D_{ref}$ and with respect to density ratios $w(x)_{ref\rightarrow prod}=\frac{p(x)_{prod}}{p(x)_{ref}}$ they become calibrated under production distribution [14]. That is true under the assumption that the true conditional probabilities of labels given the inputs remain the same for reference and production distribution, that is $p(y|X)_{ref}=p(y|X)_{prod}$. The input distribution may still change so that $p(x)_{ref}\neq p(x)_{prod}$. The probability density rates $w(x)_{ref\rightarrow prod}$ describe the covariate distribution shift from reference to production. Calibrator $c$ provides calibrated probabilities under $D_{prod}$ when:
$\mathrm{E}_{D_{ref}}\bigl{[}w(x)_{ref\rightarrow prod}\cdot y|f(x)=s\bigr{]}=c\bigl{(}f(x)\bigr{)},\forall s\in S$ (6)
@@ 3.3 Limitations
The main assumption of PAPE is that the probability distribution of labels conditioned on inputs remains the same between reference and production data, that is $p(y|X)_{ref}=p(y|X)_{prod}$. Even if that holds, the method will not work when data shifts to a region that is not covered by the reference distribution. That is when there is a region $X_{r}$ for which $p(X_{r})_{prod}>0$ but $p(X_{r})_{ref}=0$. Weighted calibration will not be possible for observations coming from that region. Another issue is that PAPE uses equation 1 that assumes that the calibrator $c$ provides calibrated probabilities, but that is rarely the case when working on finite samples of data (due to both - difficulties in learning accurate $w$ weights and the calibration process itself). Even with perfectly calibrated probabilities, the realized sample performance may end up different from the estimated one due to uncertainty related to probabilities realizing into binary targets. Even though PAPE algorithm is data-type agnostic, the DRE estimator will require different implementation for data types other than tabular data. Finally, equation 1 can be used directly to estimate a performance metric that can be defined as a mean of observation-level performance values, like accuracy. For confusion-matrix-based metrics, like F1, confusion matrix elements needs to be estimated first and then the metric of interest can be approximated. Lastly, standard concerns about using machine learning algorithms apply, such as the consequences of learning from finite samples of data.
4 Evaluation Methodology
This section describes our evaluation methodology. We describe the datasets we used in 4.1, how we created the ML models whose performance was estimated, and the experimental setup used for performance estimation evaluation in 4.2. We provide a description of the evaluated methods in 4.3. Section 4.4 shows and discusses the results. Additionally, we analyze sample size effects in 4.5.
@@ 4.1 Datasets
The datasets we use to evaluate the methods were derived with Folktables [15]. Folktables uses US Census data and preprocesses them to create a set of binary classification problems. We use the following tasks: ACSIncome, ACSPublicCoverage, ACSMobility, ACSEmployment, ACSTravelTime. For each of the five prediction tasks listed above, we fetch Folktables data for all 50 US states spanning four consecutive years (2015-2018). This gives us 250 datasets.